OpenStudy (anonymous):

4. Determine two different values of “b” in x2 + bx + 30 so that the expression can be factored into the product of two binomials. Explain how you determined those values and show each factorization. Explain how your process would change if the expression was 2x2 + bx + 30.

OpenStudy (anonymous):

b could be 17 or 13. If 17 then you have:\[(x+15)(x+2)=x^{2}+15x+2x+30\] If 13 then you have:\[(x+10)(x+3)=x^{2}+10x+3x+30\]

OpenStudy (anonymous):

And then would you just take 13 and 17 and fill it in for 'b' so it'd be... 2x^2 + (17)x + 30 and 2x^2 + (13)x + 30?

OpenStudy (anonymous):

No, that only applies to the first equation (without the leading 2). not the second one: \[x^{2}+17x+30=0\] and \[x^{2}=13x+30\]

OpenStudy (anonymous):

Oh ok. And then so for the last part I could just say that the sum will be greater because there is a ^2. right?